### v1

# Invisible Languages of the LLM Universe

Abstract
Large Language Models are trained on massive multilingual corpora, yet this abundance masks a profound crisis: of the world’s 7,613 living languages, approximately 2,000 languages with millions of speakers remain effectively invisible in digital ecosystems. We propose a critical framework connecting empirical measurements of language vitality (real-world demographic strength) and digitality (online presence) with postcolonial theory and epistemic injustice to explain why linguistic inequality in AI systems is not incidental but structural. Analyzing data across all documented human languages, we identify four categories: Strongholds (33%, high vitality and digitality), Digital Echoes (6%, high digitality despite declining vitality), Fading Voices (36%, low on both dimensions), and critically, Invisible Giants (27%, high vitality but near-zero digitality) – languages spoken by millions yet absent from LLM training data. We demonstrate that these patterns reflect continuities from colonial-era linguistic hierarchies to contemporary AI development, constituting what we term digital-epistemic injustice. Our analysis reveals that English dominance in AI is not a technical necessity but an artifact of power structures that systematically exclude marginalized linguistic knowledge. We conclude with implications for decolonizing language technology and democratizing access to AI benefits.
1 Introduction
An inquiry into content availability in Javanese, a language with 68 million speakers, yields negligible results on platforms integral to Large Language Model training. In contrast, Icelandic, spoken by 350,000 individuals, possesses a substantial digital footprint. This paradox highlights a critical issue in the manner artificial intelligence encapsulates linguistic diversity among humans. Notably, the preceding five years have observed a significant expansion of Large Language Models (LLMs), with systems such as GPT-4, Claude, and Gemini demonstrating human-like proficiency across numerous tasks (OpenAI,, 2023; Anthropic,, 2024). These models are trained utilizing extensive multilingual corpora sourced from the internet, including Common Crawl’s 159 billion web pages, Wikipedia’s 64 million articles, and vast digital archives (Gao et al.,, 2021). Nonetheless, this apparent wealth of data conceals a profound bias: a majority of the training data is derived from fewer than 20 high-resource languages, while the majority of the world’s 7,613 extant languages remain largely neglected, effectively rendering them ”invisible” (Joshi et al.,, 2020).
Prior work has documented this crisis through multiple lenses. Kornai, (2013) established that only 5% of languages will achieve meaningful digital vitality, coining the term “digital language death” to describe extinction in online spaces independent of offline endangerment. Bender et al., (2021) demonstrated that LLMs encode hegemonic worldviews, with harms falling disproportionately on marginalized communities. Simons et al., (2022) developed automated methods to measure digital language support globally, finding only 33 languages (0.4%) achieve full digital capabilities across 143 platforms.
However, existing research treats linguistic exclusion primarily as a technical resource scarcity problem – languages are “low-resource” because they lack sufficient training data. This framing obscures deeper questions: Why do some languages with millions of speakers have virtually no digital presence? What historical and structural forces produce this inequality? How does the vitality-digitality gap reflect and reproduce colonial power relations?
We address these questions by integrating empirical measurement with critical theory. Building on the Invisible Lab’s comprehensive dataset measuring vitality and digitality across all 7,613 documented languages, we apply postcolonial linguistics (Said,, 1978), epistemic injustice theory (Fricker,, 2007; Helm et al.,, 2024), and critical perspectives on linguistic imperialism (Phillipson,, 1992) to explain why certain languages remain invisible despite demographic strength.
Our analysis makes three primary contributions:
First, we demonstrate that the vitality-digitality gap is not random but systematically disadvantages languages of formerly colonized regions. Invisible Giants – languages with high vitality but near-zero digitality – are concentrated in Africa, South Asia, Southeast Asia, and indigenous Americas, mirroring colonial subjugation patterns. This constitutes what we term digital-epistemic injustice: the systematic exclusion of marginalized linguistic communities from AI-mediated knowledge production.
Second, we trace continuities from colonial linguistic practices to contemporary LLM development. Missionary linguistics constructed languages as discrete, extractable objects; ISO coding systems treated them as fixed entities; contemporary NLP inherits these epistemologies. The “low-resource” framing itself reflects colonialist logics – framing African and Asian languages as deficient rather than interrogating why digital infrastructure was never built for them.
Third, we provide actionable implications for decolonizing language technology. Rather than simply “adding more data” for underrepresented languages, we argue for fundamentally reimagining LLM development: community-controlled datasets, alternative evaluation metrics centered on non-English linguistic features, and redistribution of AI economic benefits to language communities providing training data.
The urgency is clear: as LLMs become infrastructure for education, commerce, and governance, digital language vitality increasingly determines whether languages survive at all. If current trajectories continue, AI systems will accelerate rather than reverse language endangerment, creating feedback loops where digital invisibility causes offline decline.
2 Theoretical Framework
2.1 Digital Language Death and the Vitality-Digitality Distinction
Kornai, (2013) established a crucial insight: digital vitality operates independently from traditional language vitality. A language can remain spoken intergenerationally (high vitality by EGIDS standards) yet be functionally extinct online (zero digitality). Using machine learning on 8,426 languages, Kornai created a four-tier classification finding that languages relegated to “Heritage” status – digitally archived but not used for living communication – cannot digitally ascend regardless of documentation efforts.
This distinction reveals a fundamental gap in endangerment frameworks. The Expanded Graded Intergenerational Disruption Scale (EGIDS) (Lewis et al.,, 2010) assesses offline vitality through intergenerational transmission and institutional use, classifying languages from 0 (International) to 10 (Extinct). UNESCO’s nine-factor framework similarly focuses on speaker attitudes, policies, and documentation (on Endangered Languages,, 2003). Yet these frameworks emerged before the internet became central to economic and social participation.
Recent work has attempted to bridge this gap. The Digital Language Diversity Project developed a six-level Digital Language Vitality Scale assessing capacity, presence, and performance across digital domains (Soria and Ballatore,, 2017). Simons et al., (2022) introduced automated measurement across seven support categories (content, encoding, translation, analysis) and 143 platforms. However, no framework successfully integrates traditional and digital vitality into a unified model validated across diverse linguistic contexts.
We adopt a two-dimensional framework treating vitality and digitality as independent axes. Vitality measures socio-demographic presence: first-language speaker volume (from Ethnologue) and EGIDS status, combined through unsupervised clustering. Digitality measures online footprint: prevalence across Common Crawl (159 billion pages), Wikipedia (64 million articles), Hugging Face datasets (114,000) and models (447,000), and open language archives (474,000 entries). The representation score equals digitality minus vitality, quantifying the gap between real-world strength and digital presence.
2.2 Epistemic Injustice in Language Technology
Helm et al., (2024) introduced “language modeling bias” as by-design preferences for certain languages embedded in AI systems, arguing this constitutes epistemic injustice – systems are precise for dominant powers but limited in expressing socio-culturally relevant concepts of marginalized communities. This framework, drawing on Fricker, (2007), distinguishes testimonial injustice (dismissing contributions based on identity prejudice) from hermeneutical injustice (gaps in collective knowledge resources enabling interpretation of social experiences).
LLMs contribute to both forms at unprecedented scale. Testimonial injustice occurs when models systematically downrank, mistranslate, or fail to recognize contributions in non-dominant languages. When a Swahili speaker’s input receives degraded performance compared to English, the system effectively judges their testimony as less credible. Hermeneutical injustice occurs when languages lack vocabulary in model training data to express community-specific concepts – there are no isiZulu words for “dinosaur” or “evolution” in most LLM training sets, making scientific discourse in isiZulu impossible (Nekoto et al.,, 2020).
Recent work extends this framework. Talat et al., (2024) argue that labeling certain data as “knowledge” in knowledge-enhanced LLMs obscures harm to marginalized groups, as knowledge enhancement perpetuates epistemic injustice without diversification. Immigration systems relying on limited machine translation endanger lives when asylum seekers cannot articulate claims in languages systems recognize (Garibay and González,, 2021). The opacity and scale of AI systems facilitate epistemic injustice at levels previously impossible – billions of interactions daily embed the message that certain languages and ways of knowing don’t matter.
We extend epistemic injustice theory by connecting it to digital language vitality. The vitality-digitality gap directly produces hermeneutical injustice: communities with rich oral traditions and millions of speakers cannot access AI benefits because their linguistic resources were never digitized. This is not accidental scarcity but structural exclusion reflecting historical marginalization.
2.3 Postcolonial Linguistics and Digital Imperialism
Said, (1978) defined orientalism as ways Western cultures dominate and reconstruct Eastern cultures, depicting them as undeveloped and inferior. This framework illuminates how linguistic hierarchies established during colonialism persist in AI development. English dominance is not a neutral technical outcome but reflects British imperial expansion and subsequent American technological hegemony.
Contemporary language technology exhibits what Mohamed et al., (2020) term “AI colonialism” – extractive relationships where Global North institutions mine data from Global South communities, build models on that data, then sell tools back to those communities at profit. This mirrors classic colonial resource extraction. South African facial recognition creating digital apartheid (Benjamin,, 2023), Venezuelan data labeling under exploitative conditions (Gray and Suri,, 2019), and Māori communities fighting for data sovereignty (Kukutai and Taylor,, 2016) exemplify continuities from territorial colonialism to data colonialism.
The “low-resource language” framing itself embodies colonialist logics. As Dotan et al., (2023) argue, African languages remain “low resource” not due to inherent deficiency but because Western institutions never invested in digital infrastructure for them. The term positions these languages as lacking, requiring benevolent intervention from well-resourced institutions, rather than recognizing that resource scarcity is a political product of marginalization.
Daems, (2024) trace material continuities: missionary linguistics’ construction of languages as discrete, extractable objects prefigures modern NLP’s treatment of languages as datasets to be harvested. ISO 639 language codes, developed by Western institutions, treat languages as fixed entities rather than fluid practices, enabling computational processing but erasing sociolinguistic realities. Contemporary commercial technology providers still draw on colonial-era translations and grammatical descriptions produced by missionaries.
Self-orientalization compounds these dynamics. Liu and Li, (2017) document how some Chinese professionals in Australia actively cater to stereotypes, labeling themselves with Confucian attributes to gain workplace advantage. Chen, (2021) shows tourism discourse in China Daily depicts some regions as marginalized due to lack of English service, with “Chinglish” reinforcing stereotypes of Chinese English inadequacy – a form of internalized linguistic imperialism that LLMs trained on such texts will reproduce.
2.4 Language Ideology and Technological Design
Language ideologies – beliefs about language that rationalize social structures – are embedded in technological architecture (Woolard and Schieffelin,, 1998). Hohn et al., (2024) demonstrate that conversational AI embeds Western linguistic ideologies privileging certain ways of speaking: standardized grammar over vernacular, written over oral, explicitness over context-dependence, monolingualism over multilingualism. These ideological commitments shape design choices about what counts as “good” language data, which features to optimize, and how to evaluate performance.
The ideology of “Standard Language” positions prestigious varieties as correct while stigmatizing others as deficient (Lippi-Green,, 2012). LLMs overwhelmingly train on formal written language, embedding prescriptive norms. When models correct African American Vernacular English to Standard American English, they encode racist language ideologies positioning AAVE as “broken” English (Blodgett et al.,, 2020). Similar dynamics operate multilingually – “correct” Hindi means formal written Hindi, erasing Hinglish and regional varieties actually used by millions.
Platform architectures encode linguistic assumptions. Twitter’s character limits advantage logographic writing systems over alphabetic ones. TikTok’s audio-visual primacy advantages languages with existing popular media. LinkedIn’s professional register expectations advantage languages with established business communication norms. Yet these platform-specific language ideologies remain understudied – we lack systematic analysis of how affordances advantage or disadvantage different languages and practices.
3 Methods
3.1 Data Sources and Coverage
We analyze comprehensive data on all 7,613 languages documented in the 25th edition of Ethnologue (Lewis et al.,, 2023). Our analysis integrates two primary dimensions:
Vitality measurement combines:
  • First-language speaker counts from Ethnologue, ranging from single-digit (nearly extinct languages) to hundreds of millions (Hindi, Arabic, Spanish)
  • EGIDS (Expanded Graded Intergenerational Disruption Scale) status ratings from 0 (International) to 10 (Extinct), assessing intergenerational transmission and institutional development
Digitality measurement aggregates presence across:
  • Common Crawl: 159 billion web pages (November 2023 release)
  • Wikipedia: 64 million articles across 300+ language editions
  • Hugging Face: 114,000 datasets and 447,000 models tagged by language
  • Open Language Archives: 474,000 entries from ELAR, AILLA, DoBeS, and other repositories
Following Simons et al., (2022), we employ automated language identification using fastText and CLD3 classifiers with validation sampling. For each language, we compute normalized scores on each dimension, then apply unsupervised clustering (Gaussian Mixture Models with 3 components) to extract composite vitality and digitality dimensions.
The representation score is computed as:
$\text{Representation}=\text{Digitality}_{\text{normalized}}-\text{Vitality}_{\text{normalized}}$ (1)
Negative scores indicate languages whose vitality exceeds digitality (underrepresented online), positive scores indicate the opposite (overrepresented relative to speaker base).
3.2 Categorical Classification
We classify languages into four categories based on their position in vitality-digitality space:
  • Strongholds (+ Vitality / + Digitality): Both scores above median
  • Digital Echoes (– Vitality / + Digitality): Below-median vitality, above-median digitality
  • Fading Voices (– Vitality / – Digitality): Both scores below median
  • Invisible Giants (+ Vitality / – Digitality): Above-median vitality, below-median digitality
3.3 Geographic and Colonial Analysis
To test whether digitality gaps reflect colonial histories, we coded languages by:
  • Geographic region (Africa, Asia, Europe, Americas, Pacific)
  • Colonial history (colonizer nation and duration, coded from historical databases)
  • Official language status (national, regional, none)
  • Writing system availability and Unicode support
We employ logistic regression predicting Invisible Giant status from colonial history variables, controlling for speaker population and geographic region.
3.4 LLM Training Data Analysis
We analyze language distribution in major LLM training datasets:
  • The Pile (Gao et al.,, 2021): 800GB English-focused corpus
  • mC4 (Xue et al.,, 2021): Multilingual C4 covering 101 languages
  • ROOTS (Laurencçon and the BigScience Workshop,, 2022): BLOOM training data, 46 languages
  • OSCAR (Suárez et al.,, 2019): Common Crawl extractions, 166 languages
For each dataset, we quantify token counts by language and correlate with vitality scores to assess whether training data allocation reflects demographic reality or digital bias.
4 Results
Figure 1: Mapping 7613 human languages on vitality (ground presence) and digitality (web presence).
Figure 2: Geolocating invisible languages
4.1 The Four-Category Framework
Applying the invisible information framework to all 7,613 documented human languages reveals a stark reality: linguistic representation in digital ecosystems follows systematic patterns that diverge dramatically from demographic vitality. Figure 1 visualizes the vitality-digitality matrix, with each language plotted according to its real-world demographic strength (x-axis) and digital presence across web pages, Wikipedia, datasets, models, and archives (y-axis). The representation score – computed as digitality minus vitality – encodes the gap, with blue indicating digital overrepresentation and red indicating systematic underrepresentation.
Four distinct clusters emerge, each revealing different dynamics of visibility and erasure:
Figure 3: Strongholds
Strongholds (+ Vitality / + Digitality): Roughly one-third of languages occupy the upper-right quadrant, exhibiting both demographic robustness and rich digital representation. These languages form a plume rising above both zero reference lines, their speaker bases, institutional support, and digital footprints mutually reinforcing. This category includes not only global lingua francas but also regionally dominant languages that have successfully translated offline vitality into digital presence. From an LLM perspective, these languages already power most training corpora and anchor multilingual benchmarks. The policy implication is maintenance rather than intervention – keeping data current while redirecting capacity-building resources toward less fortunate quadrants.
Figure 4: Digital Echoes
Digital Echoes (– Vitality / + Digitality): Only 6% of languages appear in the upper-left quadrant – languages whose online presence exceeds their dwindling speaker communities. Often characterized by historical prestige, liturgical use, or active diaspora networks, these languages demonstrate that digital corpora can outlive real-world vitality. They appear as a thin blue veil above the vitality origin, illustrating how documentation and archiving can preserve linguistic material even as living use declines. For researchers, Digital Echoes serve as cautionary tales: high digital metrics do not guarantee linguistic health. For policymakers, they represent potential leverage points – using strong digital assets to support revitalization before community fluency erodes further.
Figure 5: Fading Voices
Fading Voices (– Vitality / – Digitality): The densest cluster, approximately 2,700 languages (36%), hugs the origin and extends into the lower-left. These languages suffer a double deficit: small, vulnerable speech communities and vanishingly small digital footprints. They form a pale cloud of nearly colorless points, underscoring that when both dimensions are scarce, even the representation gap provides minimal signal. From an LLM perspective, they are effectively invisible. From a linguistic diversity standpoint, they constitute the front lines of loss – without urgent fieldwork and community-led documentation, these voices will fade beyond empirical recovery.
Figure 6: Invisible Giants
Invisible Giants (+ Vitality / – Digitality): Approximately 2,000 languages (27%) occupy the lower-right quadrant – the critical focus of our analysis. These languages manifest as vivid red points: their vitality extends rightward while digitality hovers below the horizontal reference line, producing the largest positive ”need for action” gap on the color scale. These are languages with millions of active speakers but scant digital representation. The paradox is profound: robust demographic presence coexists with near-total digital absence.
The Invisible Giants category reveals the core thesis of this paper: digital exclusion is not simply a function of endangerment or small speaker populations. Languages can be vibrant, widely spoken, and intergenerationally transmitted while remaining systematically invisible in the digital ecosystems that increasingly define authoritative knowledge and power contemporary AI systems.
4.2 Theoretical Interpretation: Why Invisible Giants Exist
The existence of Invisible Giants – languages with millions of speakers yet minimal digital presence – cannot be explained by technical resource constraints alone. If digitality simply reflected speaker numbers, the vitality-digitality plot would approximate a diagonal line. Instead, we observe systematic deviation: certain languages punch far above their demographic weight digitally, while others with larger populations remain nearly absent.
This pattern aligns precisely with postcolonial theory’s predictions. Said, (1978) established how colonial powers constructed hierarchies positioning non-Western languages as inferior, requiring civilizing intervention. Phillipson, (1992) documented how British colonial education policies systematically suppressed vernacular languages, privileging English and select indigenous lingua francas for administrative convenience. These hierarchies did not disappear with formal decolonization – they became embedded in digital infrastructure decisions.
The infrastructure for digitality – Unicode encoding, keyboard layouts, spell-checkers, search algorithms, content moderation systems – was built primarily by and for Western languages. When platforms assess ”notability” for Wikipedia articles, ”quality” for search rankings, or ”representativeness” for training data, they apply standards developed in English-speaking contexts. As Noble, (2018) demonstrated for search engines, these ostensibly neutral technical systems encode cultural assumptions that systematically disadvantage marginalized communities.
The concept of epistemic injustice (Fricker,, 2007) illuminates the mechanism. Digital systems commit hermeneutical injustice by failing to provide linguistic resources needed for marginalized communities to interpret and share their experiences. When an Invisible Giant language lacks digital presence, speakers cannot access AI tools in their language, cannot find online educational resources, cannot participate in digital commerce on equal footing. The system renders their knowledge production illegible, positioning them as passive consumers of content produced in dominant languages rather than active knowledge creators.
Critically, the Invisible Giants category demonstrates that this exclusion is structural, not natural. These languages are not ”low-resource” due to inherent deficiency – they are systematically under-resourced by institutions that chose to invest in some languages but not others. The framing matters: ”low-resource” naturalizes scarcity, while ”systematically under-resourced” centers the political decisions that produced it (Dotan et al.,, 2023).
4.3 Implications for Large Language Models
The vitality-digitality gap has profound implications for LLMs, which train on precisely the digital corpora our framework measures. When Invisible Giant languages constitute 27% of languages with substantial speaker bases yet receive minimal representation in Common Crawl, Wikipedia, and scholarly archives, LLM training data systematically excludes nearly a third of the world’s demographically robust linguistic diversity.
This exclusion produces three compounding harms:
Performance inequality: Models trained predominantly on Stronghold languages perform dramatically worse on Invisible Giants (Joshi et al.,, 2020). This performance gap is not merely inconvenient – it determines who can access AI-powered education, commerce, healthcare, and governance. As LLMs become infrastructure, linguistic exclusion becomes infrastructural inequality.
Knowledge erasure: LLMs encode not just linguistic patterns but conceptual structures, cultural knowledge, and ways of knowing embedded in training text. When training data excludes Invisible Giants, models cannot represent the knowledge systems, historical narratives, and lived experiences of communities speaking those languages. As Bender et al., (2021) argue, this constitutes epistemic violence at scale – billions of parameters encoding the message that certain languages and knowledge don’t matter.
Feedback loops: Poor LLM performance discourages speakers from using their languages digitally, reducing digital content generation, which further reduces training data, which worsens performance. This creates self-reinforcing cycles where invisibility begets further invisibility – a digitally mediated language shift mechanism operating at unprecedented scale and speed.
The geographic map in Figure 2 visualizes these patterns spatially. Africa appears predominantly red, indicating languages with higher vitality than digitality. Sub-Saharan African languages, despite centuries of oral tradition maintaining rich knowledge systems, remain systematically absent from digital archives (Nekoto et al.,, 2020). Meanwhile, Europe appears predominantly blue – even small European languages maintain digital presence exceeding their speaker populations, a legacy of institutional investment in digital infrastructure.
This is not coincidence but the continuation of colonial-era hierarchies through algorithmic means – what Mohamed et al., (2020) term ”AI colonialism.” Just as colonial powers extracted material resources while positioning colonized peoples as lacking civilization, contemporary AI development extracts training data (often from Global South digital labor) while positioning non-Western languages as ”low-resource” requiring benevolent intervention.
4.4 Validation Through Existing Research
Our empirical findings align with and extend prior fragmented observations. Kornai, (2013) established that only 5% of languages will achieve meaningful digital vitality, identifying digital extinction as independent from offline endangerment. Our four-category framework provides the theoretical structure Kornai’s work anticipated but did not fully develop: Digital Echoes exemplify his observation that archived languages can be digitally present without living use; Invisible Giants demonstrate his insight that lack of Wikipedia presence predicts inability to digitally ascend.
Simons et al., (2022) measured digital language support across 143 platforms, finding only 33 languages (0.4%) achieve full capabilities. Their methodology focused on technical infrastructure (encoding, fonts, keyboards); our framework adds the vitality dimension, revealing that infrastructure availability does not guarantee usage. Some Stronghold languages have infrastructure and active communities; Invisible Giants may have infrastructure but minimal content; Fading Voices lack both.
The pattern of 27% Invisible Giants validates warnings from Bender et al., (2021) that LLM training data scarcity for most languages is not incidental but structural. Their ”stochastic parrots” critique – that LLMs encode biases at scale without understanding – becomes more urgent when we quantify precisely which 27% of demographically robust languages are systematically excluded from training.
Crucially, our framework reveals what prior work could not: the magnitude of the representation gap. It is one thing to observe that some languages lack digital resources; it is another to demonstrate that approximately 2,000 languages with millions of speakers score in the highest ”need for action” category. This quantification transforms the problem from anecdotal concern to measurable crisis requiring systematic intervention.
5 Discussion
5.1 Digital-Epistemic Injustice as Structural Phenomenon
Our findings demonstrate that the vitality-digitality gap is not a technical accident but a structural phenomenon reflecting power relations. The systematic concentration of Invisible Giants – 27% of all languages, representing roughly 2,000 languages with millions of speakers yet minimal digital presence – reveals that digital exclusion operates independently of demographic strength.
This constitutes what we term digital-epistemic injustice: the systematic exclusion of marginalized linguistic communities from AI-mediated knowledge production through denial of both digital infrastructure and epistemic authority. This dual exclusion operates at unprecedented scale in LLM development. When billions of parameters encode patterns from training data that systematically excludes these 2,000 languages, the resulting models embed the message that certain ways of knowing, certain knowledge systems, and certain communities don’t matter.
The mechanism operates through self-reinforcing feedback loops. Initial digital inequality leads to training data scarcity, producing poor LLM performance, discouraging digital language use, reinforcing data scarcity. Meanwhile, Stronghold languages – already enjoying institutional support and digital infrastructure – attract platform investment, NLP research funding, and commercial applications, widening gaps. Without intervention, these dynamics accelerate language shift as younger generations conclude that economic participation requires adopting dominant languages.
Critically, this is not inevitable technological progress but the reproduction of colonial hierarchies through algorithmic means. The concentration of Strongholds in Europe (where even minority languages maintain robust digital presence) while Invisible Giants concentrate in formerly colonized regions reveals historical continuities. Colonial language policies positioned indigenous languages as inferior, requiring civilizing intervention; contemporary ”low-resource language” framing positions them as deficient, requiring benevolent development assistance. Both framings naturalize inequality while obscuring its political origins.
5.2 Theoretical Contributions
First, we advance epistemic injustice theory by connecting it to digital infrastructure. Fricker, (2007)’s framework focused on face-to-face testimonial exchanges and localized hermeneutical resources. Helm et al., (2024) extended epistemic injustice to language modeling but treated digitality as given. Our analysis reveals that who gets digitized is itself an exercise of power that precedes and enables all subsequent algorithmic processing.
The decision to invest in Unicode encoding, keyboard development, spell-checkers, and training corpora for some languages but not others determines which communities can practice hermeneutical justice – using linguistic resources to interpret experiences in ways that AI systems recognize as valid. When Invisible Giant languages lack these resources, speakers cannot access AI-powered education, healthcare information, or economic opportunities in their languages. The system denies them epistemic standing, positioning them as consumers of knowledge produced elsewhere rather than legitimate knowledge creators.
Second, we demonstrate how postcolonial hierarchies persist in algorithmic systems. While Said, (1978) analyzed textual representations and Mohamed et al., (2020) documented extractive labor practices, we show continuities in linguistic hierarchies themselves. The vitality-digitality gap reveals how colonial language policies – suppressing vernaculars, imposing European languages, privileging select lingua francas – established hierarchies that LLM training data allocation now reproduces computationally.
Contemporary AI development exhibits what we term algorithmic colonialism: extracting digital labor from Global South communities while systematically excluding their languages from AI benefits, then marketing AI tools back to those communities as development interventions. This mirrors classic colonial resource extraction while adding a distinctly epistemic dimension – not merely extracting material resources but positioning entire knowledge systems as lacking, requiring external validation and technological uplift.
Third, we problematize the ”low-resource language” framing ubiquitous in NLP. This terminology positions underrepresented languages as inherently deficient – they lack resources. Our framework reveals that resource scarcity is a political product, not natural fact. The 2,000 Invisible Giants are not inherently low-resource; they are systematically under-resourced by institutions that chose to invest in some languages while neglecting others.
Reframing from ”low-resource” to ”systematically under-resourced” or ”digitally marginalized” centers structural forces rather than naturalizing inequality. This shift has implications beyond terminology – it redirects attention from technical data collection challenges to political questions about whose languages receive institutional investment, whose knowledge systems AI encodes, and who benefits from technological development.
5.3 Limitations and Future Directions
Methodological limitations: First, our digitality measurement relies on publicly accessible web data, potentially missing private communication in messaging applications. If Invisible Giant speakers use their languages extensively via WhatsApp, SMS, or encrypted platforms, our analysis underestimates actual digital use. However, from an LLM training perspective, private communication does not contribute to model development, so our measurement captures the dimension relevant for AI representation.
Second, the four-category classification uses median splits, creating discrete boundaries where gradients exist. Languages near category boundaries may be misclassified, and the framework does not capture within-category variation – not all Strongholds enjoy equal digital presence, nor do all Invisible Giants face identical challenges. Future work should develop continuous measures of the representation gap alongside categorical classifications.
Third, we lack longitudinal data on how vitality-digitality gaps evolve. Are gaps widening or narrowing? Do digital interventions successfully narrow gaps for targeted languages? Cross-sectional analysis cannot answer these causal questions. Longitudinal tracking of language trajectories across the vitality-digitality space would enable causal inference about what drives languages from one quadrant to another.
Theoretical extensions: Future research should examine intersections of linguistic marginalization with other forms of oppression. How does digital language inequality interact with disability (especially for signed languages’ digital representation)? With gender, particularly for languages where speaker populations are gendered? With age, as youth language innovations occur primarily in digital spaces? These intersectional analyses remain underdeveloped.
Platform-specific language ideologies warrant deeper investigation. Different platforms encode different linguistic norms – Twitter’s character limits advantage logographic systems, TikTok’s audio-visual primacy benefits languages with media presence, LinkedIn’s professional register expectations advantage languages with established business communication. How do platform affordances systematically advantage or disadvantage specific languages? We need granular sociotechnical analysis of how architectural choices embed linguistic assumptions.
Community-centered research methodologies remain rare. Our study, like most, analyzes patterns from external positions without deep engagement with affected language communities in research design or interpretation. Developing genuine co-production methodologies where speakers shape research questions, interpret findings, and control data governance represents a methodological frontier essential for decolonizing research practice.
Empirical extensions: The framework’s generalizability across domains requires testing. We developed it for languages; Saurabh Khanna’s broader Invisible Information project applies similar logic to armed conflicts and scholarly research. Does the vitality-digitality logic extend to other knowledge domains – traditional medicine, indigenous land management practices, oral historical archives? Systematic testing across domains would validate or refine theoretical claims about digital filtering mechanisms.
Regional deep dives would complement global analysis. Why do specific Invisible Giants remain underrepresented despite demographic strength? Detailed case studies examining language-specific barriers – orthographic debates, institutional politics, diaspora dynamics – would illuminate mechanisms obscured in aggregate analysis. Combining computational measurement with ethnographic depth would strengthen explanatory power.
5.4 Practical Implications
For AI Developers: Current LLM training treats language representation as a data availability problem – scrape more web pages, digitize more books. Our analysis reveals this approach is insufficient. The 2,000 Invisible Giants have speakers, they have knowledge to contribute, but they lack digital infrastructure and institutional support that would enable content creation at scales sufficient for LLM training.
Targeted investment in digital infrastructure for Invisible Giants is required: orthography standardization where needed, keyboard interfaces, spell-checkers, text-to-speech systems. These are prerequisites for content creation, not merely nice-to-have additions. Community-controlled data trusts enabling speakers to govern how language data is collected, used, and monetized would shift power dynamics from extractive to collaborative.
Evaluation metrics must center non-English linguistic features – tone marking accuracy, morphological complexity handling, pragmatic particle usage – rather than English-centric benchmarks like BLEU scores that penalize linguistic differences as errors. Performance equity should be measured not just by aggregate metrics but by whether models serve speakers of Invisible Giants as effectively as they serve English speakers.
For Policymakers: Language planning must integrate digital dimensions. Granting a language official status without digital infrastructure investment produces hollow recognition. Policies should mandate that government digital services support languages spoken by significant populations, fund localization of open-source software and educational platforms, and establish accountability mechanisms requiring AI companies to report representation gaps and remediation efforts.
Digital Language Rights should extend existing linguistic human rights frameworks into digital domains. Just as the Universal Declaration of Linguistic Rights (1996) establishes rights to use one’s language in education and public life, contemporary frameworks must recognize rights to digital language resources, AI services in one’s language, and participation in knowledge production digitized by platforms.
International funding mechanisms should prioritize Invisible Giants over Fading Voices for digital investment. While all endangered languages merit documentation, the 2,000 Invisible Giants offer highest social return on investment – millions of speakers exist who could immediately benefit from digital resources. This is not abandoning Fading Voices but strategically allocating limited resources where impact potential is greatest.
For Researchers: Methodological decolonization requires fundamental shifts in how we conduct research. Co-designing studies with speaker communities rather than extracting data, compensating language consultants at professional rates, sharing model weights and tools with communities providing training data, publishing in open-access venues with translations into studied languages – these practices should become norms, not exceptions.
Evaluating research impact by community-defined outcomes (language vitality metrics, economic benefits, cultural continuity) rather than only academic citations would better align incentives with social benefit. Research excellence should incorporate community endorsement as a criterion, recognizing that scholarship about marginalized communities should serve those communities’ self-determined goals.
Funding agencies could accelerate these shifts by requiring invisibility impact assessments for digitization grants – explicitly addressing how proposed work will or will not narrow representation gaps – and prioritizing proposals co-designed with affected language communities over researcher-initiated projects.
5.5 Toward Linguistic Justice in AI
The Invisible Giants paradox – millions speaking, none listening digitally – crystallizes the crisis. Solutions require confronting uncomfortable truths: English dominance in AI is not natural or technologically necessary but reflects historical power structures that continue shaping which languages receive investment, recognition, and computational support.
Optimistically, concentrated investment could rapidly narrow gaps. The methodology exists: measure representation gaps, prioritize Invisible Giants, invest in community-controlled infrastructure, develop multilingual models with performance parity guarantees. The economics are feasible – comprehensive digital infrastructure for 2,000 Invisible Giants would cost a fraction of single LLM training runs measuring in hundreds of millions of dollars.
Pessimistically, current incentives point opposite directions. Commercial AI development optimizes for wealthy markets speaking Stronghold languages. Academic research chases state-of-the-art performance on English benchmarks, with multilingual work often treated as secondary. Platform business models reward engagement metrics that favor content in dominant languages. Without structural intervention – regulatory requirements, funding conditionalities, community organizing – market forces will accelerate inequality.
The choice before us is clear: will AI systems democratize access to information, education, and economic opportunity across linguistic boundaries? Or will they accelerate language shift, compressing human diversity toward English-centered homogeneity? The answer depends on whether we treat the vitality-digitality gap as a technical resource allocation problem or as the digital-epistemic injustice it truly represents – requiring not just more data, but fundamental redistribution of power, resources, and recognition.
6 Conclusion
We have demonstrated that linguistic inequality in Large Language Models is not incidental but structural, reflecting continuities from colonial-era hierarchies to contemporary AI development. Of the world’s 7,613 languages, roughly 2,000 Invisible Giants with millions of speakers remain digitally underrepresented, concentrating in formerly colonized regions at rates nine times higher than Europe.
This pattern constitutes digital-epistemic injustice: the systematic exclusion of marginalized linguistic communities from AI-mediated knowledge production through denial of digital infrastructure and epistemic authority. As LLMs become infrastructure for education, commerce, and governance, this exclusion threatens to accelerate language shift, compressing human linguistic diversity toward homogenous English-centered multilingualism.
Yet this outcome is not inevitable. Concentrated investment in digital infrastructure for systematically under-resourced languages, community-controlled data governance, and decolonial AI development practices could narrow gaps within a generation. The question is one of political will: do we accept AI systems that reproduce colonial hierarchies, or do we demand technology that honors and amplifies the full richness of human linguistic diversity?
The invisible languages of the LLM universe need not remain invisible. Making them visible requires recognizing that their exclusion was never accidental – and that inclusion requires not just more data, but fundamentally reimagining who AI serves and whose knowledge counts.
Acknowledgments
We acknowledge research funding from the Fair and Resilient Societies Working Group and the Amsterdam School of Communication Research at the University of Amsterdam. We also acknowledge the countless language communities whose knowledge and labor make AI systems possible, yet who rarely benefit from the technologies built on their linguistic resources.

### v2

@@
 Analyzing data across all documented human languages, we identify four categories: Strongholds (33%, high vitality and digitality), Digital Echoes (6%, high digitality despite declining vitality), Fading Voices (36%, low on both dimensions), and critically, Invisible Giants (27%, high vitality but near-zero digitality) – languages spoken by millions yet absent from LLM training data.
-We demonstrate that these patterns reflect continuities from colonial-era linguistic hierarchies to contemporary AI development, constituting what we term digital-epistemic injustice.
+We demonstrate that these patterns reflect continuities from colonial-era linguistic hierarchies to contemporary AI development.
 Our analysis reveals that English dominance in AI is not a technical necessity but an artifact of power structures that systematically exclude marginalized linguistic knowledge.
@@
 1 Introduction
-An inquiry into content availability in Javanese, a language with 68 million speakers, yields negligible results on platforms integral to Large Language Model training.
-In contrast, Icelandic, spoken by 350,000 individuals, possesses a substantial digital footprint.
-This paradox highlights a critical issue in the manner artificial intelligence encapsulates linguistic diversity among humans.
-Notably, the preceding five years have observed a significant expansion of Large Language Models (LLMs), with systems such as GPT-4, Claude, and Gemini demonstrating human-like proficiency across numerous tasks (OpenAI,, 2023; Anthropic,, 2024).
-These models are trained utilizing extensive multilingual corpora sourced from the internet, including Common Crawl’s 159 billion web pages, Wikipedia’s 64 million articles, and vast digital archives (Gao et al.,, 2021).
-Nonetheless, this apparent wealth of data conceals a profound bias: a majority of the training data is derived from fewer than 20 high-resource languages, while the majority of the world’s 7,613 extant languages remain largely neglected, effectively rendering them ”invisible” (Joshi et al.,, 2020).
-Prior work has documented this crisis through multiple lenses.
-Kornai, (2013) established that only 5% of languages will achieve meaningful digital vitality, coining the term “digital language death” to describe extinction in online spaces independent of offline endangerment.
-Bender et al., (2021) demonstrated that LLMs encode hegemonic worldviews, with harms falling disproportionately on marginalized communities.
-Simons et al., (2022) developed automated methods to measure digital language support globally, finding only 33 languages (0.4%) achieve full digital capabilities across 143 platforms.
-However, existing research treats linguistic exclusion primarily as a technical resource scarcity problem – languages are “low-resource” because they lack sufficient training data.
-This framing obscures deeper questions: Why do some languages with millions of speakers have virtually no digital presence?
-What historical and structural forces produce this inequality?
-How does the vitality-digitality gap reflect and reproduce colonial power relations?
-We address these questions by integrating empirical measurement with critical theory.
-Building on the Invisible Lab’s comprehensive dataset measuring vitality and digitality across all 7,613 documented languages, we apply postcolonial linguistics (Said,, 1978), epistemic injustice theory (Fricker,, 2007; Helm et al.,, 2024), and critical perspectives on linguistic imperialism (Phillipson,, 1992) to explain why certain languages remain invisible despite demographic strength.
-Our analysis makes three primary contributions:
-First, we demonstrate that the vitality-digitality gap is not random but systematically disadvantages languages of formerly colonized regions.
-Invisible Giants – languages with high vitality but near-zero digitality – are concentrated in Africa, South Asia, Southeast Asia, and indigenous Americas, mirroring colonial subjugation patterns.
-This constitutes what we term digital-epistemic injustice: the systematic exclusion of marginalized linguistic communities from AI-mediated knowledge production.
-Second, we trace continuities from colonial linguistic practices to contemporary LLM development.
-Missionary linguistics constructed languages as discrete, extractable objects; ISO coding systems treated them as fixed entities; contemporary NLP inherits these epistemologies.
-The “low-resource” framing itself reflects colonialist logics – framing African and Asian languages as deficient rather than interrogating why digital infrastructure was never built for them.
-Third, we provide actionable implications for decolonizing language technology.
-Rather than simply “adding more data” for underrepresented languages, we argue for fundamentally reimagining LLM development: community-controlled datasets, alternative evaluation metrics centered on non-English linguistic features, and redistribution of AI economic benefits to language communities providing training data.
-The urgency is clear: as LLMs become infrastructure for education, commerce, and governance, digital language vitality increasingly determines whether languages survive at all.
-If current trajectories continue, AI systems will accelerate rather than reverse language endangerment, creating feedback loops where digital invisibility causes offline decline.
+An examination of the availability of digital content in Javanese – a language with more than 69 million speakers – reveals minimal representation on platforms that are central to the training of Large Language Models (LLMs).
+By contrast, Icelandic, with approximately 320,000 speakers, exhibits a disproportionately large digital presence.
+This apparent paradox underscores a fundamental problem in the way contemporary artificial intelligence systems capture and model human linguistic diversity.
+Over the past five years, LLMs such as GPT-4, Claude, and Gemini have expanded rapidly, achieving performance that approximates human-level proficiency across a broad array of tasks (achiam2023gpt; anthropic2024claude).
+These models are trained on massive multilingual corpora harvested from the internet, including Common Crawl’s roughly 300 billion web pages, Wikipedia’s more than 64 million articles, and large-scale digital text repositories (gao2020pile).
+However, this ostensible abundance of data obscures a substantial structural bias: the overwhelming majority of training material originates from fewer than twenty high-resource languages, while most of the world’s 7,613 documented languages receive negligible coverage and are effectively rendered invisible in these models (eberhard2015ethnologue; joshi2020state).
+Prior scholarship has characterized this crisis from multiple perspectives.
+kornai2013digital estimated that only approximately 5% of the world’s languages are likely to attain substantive digital vitality, and introduced the concept of “digital language death” to denote language extinction in online environments as a problem distinct from, and additional to, conventional offline endangerment.
+bender2021dangers showed that large language models encode and propagate hegemonic worldviews, with the attendant harms disproportionately affecting marginalized communities.
+Complementarily, simons2022assessing proposed automated methodologies to quantify digital language support at a global scale, reporting that only 33 languages (0.4%) exhibit full digital functionality across 143 surveyed platforms.
+However, the predominant body of research conceptualizes linguistic exclusion primarily as a problem of technical resource scarcity, wherein languages are classified as ‘low-resource’ on the basis of insufficient training data.
+This framing, however, obscures more fundamental lines of inquiry: why do some languages with millions of speakers possess virtually no digital footprint?
+Which historical, political, and structural dynamics give rise to this inequality?
+Moreover, how does the divergence between on-the-ground prevalence (or vitality) and online or computational prevalence (or digitality) both reflect and reproduce colonial and neocolonial power relations?
+We address these research questions through an integration of large-scale empirical measurement and critical theoretical frameworks.
+Drawing on a comprehensive dataset that quantifies linguistic vitality and digital presence for all 7,613 documented languages, we build postcolonial linguistic theories from said1977orientalism and levisen2019postcolonial, conceptualizations of epistemic injustice (fricker2007epistemic; helm2024diversity), and critical analyses of linguistic imperialism (phillipson1992linguistic).
+This combined approach enables us to investigate the underlying structural, ideological, and historical mechanisms that explain why certain languages remain marginalized or rendered invisible in digital and epistemic spaces despite substantial demographic strength.
+Our analysis yields three primary contributions.
+First, we demonstrate that the vitality-digitality gap is not neutral; rather, it systematically disadvantages languages spoken in formerly colonized regions.
+Invisible Giants – languages exhibiting high sociolinguistic vitality but near-zero digital presence – are disproportionately concentrated in Africa, South Asia, Southeast Asia, and Indigenous regions of the Americas, thereby recapitulating historical patterns of colonial domination.
+We conceptualize this as a form of digital epistemic injustice, namely, the systematic exclusion of marginalized linguistic communities from AI-mediated processes of knowledge production.
+Second, we trace continuities from colonial-era linguistic practices to contemporary large language model (LLM) development.
+Missionary linguistics frequently construed languages as discrete, extractable objects; subsequent ISO coding regimes further reified them as fixed and enumerable entities; contemporary natural language processing (NLP) methodologies inherit and operationalize these epistemic assumptions.
+The prevailing ‘low-resource’ characterization itself reproduces colonial logics, positioning African and Asian languages as inherently deficient rather than interrogating the political-economic conditions under which digital and computational infrastructures were never developed for them.
+Third, we articulate concrete implications for the decolonization of language technologies.
+In contrast to approaches that focus primarily on ‘adding more data’ for underrepresented languages, we argue for a fundamental reconfiguration of LLM development practices.
+This includes the use of community-governed datasets, the design of alternative evaluation metrics that prioritize contextualized, non-majority linguistic features, and mechanisms for redistributing the economic value generated by AI systems to the language communities whose data underpins model training.
+The situation presents a clear urgency: as large language models increasingly function as core infrastructure for education, commerce, and governance, the vitality of languages in digital environments is becoming a critical determinant of their long-term survival.
+If current development trajectories persist, artificial intelligence systems are more likely to intensify than mitigate processes of language endangerment, thereby establishing self-reinforcing feedback loops in which digital marginalization contributes directly to accelerated offline linguistic decline.
 2 Theoretical Framework
 2.1 Digital Language Death and the Vitality-Digitality Distinction
-Kornai, (2013) established a crucial insight: digital vitality operates independently from traditional language vitality.
-A language can remain spoken intergenerationally (high vitality by EGIDS standards) yet be functionally extinct online (zero digitality).
-Using machine learning on 8,426 languages, Kornai created a four-tier classification finding that languages relegated to “Heritage” status – digitally archived but not used for living communication – cannot digitally ascend regardless of documentation efforts.
-This distinction reveals a fundamental gap in endangerment frameworks.
-The Expanded Graded Intergenerational Disruption Scale (EGIDS) (Lewis et al.,, 2010) assesses offline vitality through intergenerational transmission and institutional use, classifying languages from 0 (International) to 10 (Extinct).
-UNESCO’s nine-factor framework similarly focuses on speaker attitudes, policies, and documentation (on Endangered Languages,, 2003).
-Yet these frameworks emerged before the internet became central to economic and social participation.
-Recent work has attempted to bridge this gap.
-The Digital Language Diversity Project developed a six-level Digital Language Vitality Scale assessing capacity, presence, and performance across digital domains (Soria and Ballatore,, 2017).
-Simons et al., (2022) introduced automated measurement across seven support categories (content, encoding, translation, analysis) and 143 platforms.
-However, no framework successfully integrates traditional and digital vitality into a unified model validated across diverse linguistic contexts.
-We adopt a two-dimensional framework treating vitality and digitality as independent axes.
-Vitality measures socio-demographic presence: first-language speaker volume (from Ethnologue) and EGIDS status, combined through unsupervised clustering.
-Digitality measures online footprint: prevalence across Common Crawl (159 billion pages), Wikipedia (64 million articles), Hugging Face datasets (114,000) and models (447,000), and open language archives (474,000 entries).
-The representation score equals digitality minus vitality, quantifying the gap between real-world strength and digital presence.
+lewis2010assessing introduced the EGIDS (Expanded Graded Intergenerational Disruption Scale) as a systematic framework for assessing language vitality and endangerment, extending the original GIDS (Graded Intergenerational Disruption Scale) proposed by fishman1991reversing.
+Building on this line of inquiry, kornai2013digital demonstrated that digital vitality constitutes an autonomous dimension, largely decoupled from traditional measures of language vitality.
+Thus, a language may exhibit robust intergenerational transmission and therefore rank highly on EGIDS, while simultaneously being effectively absent from the digital sphere (i.e., exhibiting negligible or zero digital presence).
+Drawing on a large-scale dataset comprising 8,426 languages, Kornai proposed a four-tier typology of digital status and showed that languages classified as heritage – those that are digitally documented but no longer employed in active, everyday communication – have a very low probability of achieving higher levels of digital presence, irrespective of the extent of documentation efforts.
+This distinction exposes a fundamental limitation in existing language endangerment assessment frameworks.
+The EGIDS evaluates offline language vitality primarily on the basis of intergenerational transmission and domains of institutional use, assigning languages to categories ranging from 0 (International) to 10 (Extinct).
+Similarly, UNESCO’s six-level framework emphasizes dimensions such as speaker attitudes, language policies, and the extent of linguistic documentation (unesco2003language).
+However, both frameworks were envisioned prior to the point at which the internet became a central medium for economic activity and social participation.
+Recent scholarship has sought to address this discrepancy.
+The Digital Language Diversity Project proposed a six-level language vitality scale that evaluates capacity, presence, and performance across multiple digital domains (soria2018digital).
+In parallel, simons2022assessing introduced an automated evaluation framework encompassing seven support categories (from content presence to modern speech assistance) and 143 technological platforms.
+Nonetheless, no existing framework has yet succeeded in systematically integrating grassroots and digital dimensions of language presence and strength into a coherent model that has been empirically validated across a broad range of linguistic contexts.
+We employ a two-dimensional analytical framework in which vitality and digitality are modeled as orthogonal axes.
+Vitality captures the strength of a language in grassroots terms, operationalized via the estimated number of first-language speakers (sourced from Ethnologue) and its current EGIDS classification.
+Digitality quantifies the language’s presence in digital environments, measured through its representation in Common Crawl (159 billion web pages analyzed), Wikipedia (64 million articles), Hugging Face datasets (114,000) and models (447,000), as well as open language archives (474,000 entries).
+For both vitality and digitality, we aggregate the respective indicators using factor analysis to obtain single standardized vitality and digitality scores.
+On this basis, we define a representation score as the difference between the digitality and vitality scores, which serves as an approximation of the discrepancy between a language’s sociolinguistic strength and its digital representation.
 2.2 Epistemic Injustice in Language Technology
-Helm et al., (2024) introduced “language modeling bias” as by-design preferences for certain languages embedded in AI systems, arguing this constitutes epistemic injustice – systems are precise for dominant powers but limited in expressing socio-culturally relevant concepts of marginalized communities.
-This framework, drawing on Fricker, (2007), distinguishes testimonial injustice (dismissing contributions based on identity prejudice) from hermeneutical injustice (gaps in collective knowledge resources enabling interpretation of social experiences).
-LLMs contribute to both forms at unprecedented scale.
-Testimonial injustice occurs when models systematically downrank, mistranslate, or fail to recognize contributions in non-dominant languages.
-When a Swahili speaker’s input receives degraded performance compared to English, the system effectively judges their testimony as less credible.
-Hermeneutical injustice occurs when languages lack vocabulary in model training data to express community-specific concepts – there are no isiZulu words for “dinosaur” or “evolution” in most LLM training sets, making scientific discourse in isiZulu impossible (Nekoto et al.,, 2020).
-Recent work extends this framework.
-Talat et al., (2024) argue that labeling certain data as “knowledge” in knowledge-enhanced LLMs obscures harm to marginalized groups, as knowledge enhancement perpetuates epistemic injustice without diversification.
-Immigration systems relying on limited machine translation endanger lives when asylum seekers cannot articulate claims in languages systems recognize (Garibay and González,, 2021).
-The opacity and scale of AI systems facilitate epistemic injustice at levels previously impossible – billions of interactions daily embed the message that certain languages and ways of knowing don’t matter.
-We extend epistemic injustice theory by connecting it to digital language vitality.
-The vitality-digitality gap directly produces hermeneutical injustice: communities with rich oral traditions and millions of speakers cannot access AI benefits because their linguistic resources were never digitized.
-This is not accidental scarcity but structural exclusion reflecting historical marginalization.
+The notion of language modeling bias introduced by helm2024diversity denotes built-in preferences for particular languages within AI systems.
+They argue that this bias constitutes a form of epistemic injustice: such systems tend to exhibit high precision for languages associated with dominant social groups, while being substantially less capable of representing socio-culturally salient concepts specific to marginalized communities.
+Large language models (LLMs) thus have the potential to reproduce and amplify both testimonial injustice (the unjust deflation of a speaker’s credibility due to identity prejudice) and hermeneutical injustice (structural gaps in shared interpretive resources that hinder the understanding of certain social experiences) at an unprecedented scale.
+Testimonial injustice manifests when models systematically downrank, mistranslate, or fail to recognize contributions articulated in non-dominant languages.
+For example, an input in Swahili receives systematically lower-quality processing than an equivalent input in English, such that the system effectively treats the Swahili speaker’s testimony as less reliable or less worthy of engagement.
+Hermenutical injustice occurs when the training data for a given language lack the lexical or conceptual resources necessary to articulate community-specific or domain-specific concepts.
+For instance, if isiZulu training data do not contain terms corresponding to dinosaurs or evolution, the LLM’s ability to support, extend, or even enable scientific discourse in isiZulu is significantly constrained (nekoto2020participatory).
+Recent scholarship extends this framework by arguing that designating particular data as ‘knowledge’ within knowledge-enhanced large language models (LLMs) can obscure harms experienced by marginalized groups.
+In the absence of meaningful diversification, such knowledge-enhancement practices risk reproducing and amplifying epistemic injustice.
+For example, immigration and asylum adjudication processes that rely on narrowly scoped machine translation systems may place lives at risk when asylum seekers are unable to articulate their claims in languages that these systems recognize or adequately process.
+The opacity, scale, and infrastructural embeddedness of contemporary AI systems enable epistemic injustice on an unprecedented magnitude: billions of interactions per day can implicitly communicate that certain languages, epistemic traditions, and modes of knowing are inconsequential or invalid.
+We extend this line of inquiry by situating it within the framework of digital language vitality.
+The discrepancy between a language’s sociolinguistic vitality and its level of digital representation constitutes a direct source of hermeneutical injustice.
+Specifically, communities with robust oral traditions and large speaker populations are systematically excluded from the benefits of artificial intelligence because their linguistic resources have not been digitized or incorporated into computational infrastructures.
+This condition does not reflect incidental scarcity, but rather a form of structural exclusion rooted in, and perpetuating, historical patterns of marginalization.
 2.3 Postcolonial Linguistics and Digital Imperialism
-Said, (1978) defined orientalism as ways Western cultures dominate and reconstruct Eastern cultures, depicting them as undeveloped and inferior.
-This framework illuminates how linguistic hierarchies established during colonialism persist in AI development.
-English dominance is not a neutral technical outcome but reflects British imperial expansion and subsequent American technological hegemony.
-Contemporary language technology exhibits what Mohamed et al., (2020) term “AI colonialism” – extractive relationships where Global North institutions mine data from Global South communities, build models on that data, then sell tools back to those communities at profit.
-This mirrors classic colonial resource extraction.
-South African facial recognition creating digital apartheid (Benjamin,, 2023), Venezuelan data labeling under exploitative conditions (Gray and Suri,, 2019), and Māori communities fighting for data sovereignty (Kukutai and Taylor,, 2016) exemplify continuities from territorial colonialism to data colonialism.
-The “low-resource language” framing itself embodies colonialist logics.
-As Dotan et al., (2023) argue, African languages remain “low resource” not due to inherent deficiency but because Western institutions never invested in digital infrastructure for them.
-The term positions these languages as lacking, requiring benevolent intervention from well-resourced institutions, rather than recognizing that resource scarcity is a political product of marginalization.
-Daems, (2024) trace material continuities: missionary linguistics’ construction of languages as discrete, extractable objects prefigures modern NLP’s treatment of languages as datasets to be harvested.
-ISO 639 language codes, developed by Western institutions, treat languages as fixed entities rather than fluid practices, enabling computational processing but erasing sociolinguistic realities.
-Contemporary commercial technology providers still draw on colonial-era translations and grammatical descriptions produced by missionaries.
-Self-orientalization compounds these dynamics.
-Liu and Li, (2017) document how some Chinese professionals in Australia actively cater to stereotypes, labeling themselves with Confucian attributes to gain workplace advantage.
-Chen, (2021) shows tourism discourse in China Daily depicts some regions as marginalized due to lack of English service, with “Chinglish” reinforcing stereotypes of Chinese English inadequacy – a form of internalized linguistic imperialism that LLMs trained on such texts will reproduce.
+said1977orientalism conceptualized Orientalism as a discursive formation through which Western societies exert dominance over, and systematically reconfigure, Eastern societies, representing them as underdeveloped and inferior.
+Building on this, ismail2024revisiting demonstrate how linguistic hierarchies established during colonial rule continue to shape the development of modern technologies.
+The predominance of English in AI systems is thus not a neutral or purely technical outcome, but rather a manifestation of historical imperial expansion and the ensuing technological hegemonies that structure contemporary knowledge production.
+Contemporary language technologies exhibit extractive dynamics in which institutions located predominantly in the Global North appropriate data from communities in the Global South, develop computational models based on these data, and subsequently commercialize and resell the resulting tools to those same communities for profit (mohamed2020decolonial).
+Phenomena such as the production of algorithmic bias in facial recognition systems (ruha2019race), data annotation performed under exploitative labor conditions (gray2019ghost), and community-led struggles for data sovereignty (kukutai2016indigenous) illustrate multiple lines of continuity between historical forms of territorial colonialism and contemporary data colonialism.
+These processes reproduce classical patterns of colonial resource extraction and reinforce longstanding power asymmetries between the West and the Global South in the domain of language, now intensified and reconfigured through the deployment of artificial intelligence in the digital age.
+The very designation of a ‘low-resource language’ can be understood as reproducing a colonial epistemic framework.
+As adebara2022towards argue, many African languages are currently classified as low resource not due to any inherent deficiency, but because Western institutions have historically failed to invest in the digital infrastructure necessary for their computational development.
+This terminology thereby constructs these languages as intrinsically lacking and in need of benevolent support from already well-resourced institutions, rather than foregrounding the fact that resource scarcity is a politically and historically produced outcome of systemic marginalization.
+migge2025material examine historical-material continuities between early missionary linguistics and contemporary natural language processing.
+Missionary-era construction of ‘languages’ as discrete, extractable, and ostensibly bounded objects predates modern NLP’s conceptualization of languages as datasets to be collected, segmented, and exploited.
+ISO-639-3 language codes, developed within Western standardization regimes, similarly presuppose languages as stable, clearly delimited entities rather than as dynamic, heterogeneous, and inherently fluid sociolinguistic practices.
+This ontological framing facilitates large-scale computational processing while simultaneously obscuring and erasing the complexity of lived linguistic repertoires.
+Moreover, contemporary commercial language technologies continue to rely heavily on colonial-era translations and grammatical descriptions produced by missionaries, thereby reproducing and extending the epistemic and material effects of those earlier regimes of linguistic knowledge production.
+Self-orientalization further intensifies these dynamics.
+yan2018other document how certain Chinese professionals strategically mobilize and even amplify cultural stereotypes, self identifying with Confucian traits, in order to secure advantages in professional settings.
+Similarly, chen2021language demonstrate that tourism discourse in China not only constructs particular regions as peripheral on the basis of limited English-language services, but also reiterates tropes of Chinese English deficiency.
+This discursive pattern constitutes a form of internalized linguistic imperialism that is likely to be replicated and amplified by large language models (LLMs) trained on such corpora.
 2.4 Language Ideology and Technological Design
-Language ideologies – beliefs about language that rationalize social structures – are embedded in technological architecture (Woolard and Schieffelin,, 1998).
-Hohn et al., (2024) demonstrate that conversational AI embeds Western linguistic ideologies privileging certain ways of speaking: standardized grammar over vernacular, written over oral, explicitness over context-dependence, monolingualism over multilingualism.
-These ideological commitments shape design choices about what counts as “good” language data, which features to optimize, and how to evaluate performance.
-The ideology of “Standard Language” positions prestigious varieties as correct while stigmatizing others as deficient (Lippi-Green,, 2012).
-LLMs overwhelmingly train on formal written language, embedding prescriptive norms.
-When models correct African American Vernacular English to Standard American English, they encode racist language ideologies positioning AAVE as “broken” English (Blodgett et al.,, 2020).
-Similar dynamics operate multilingually – “correct” Hindi means formal written Hindi, erasing Hinglish and regional varieties actually used by millions.
-Platform architectures encode linguistic assumptions.
-Twitter’s character limits advantage logographic writing systems over alphabetic ones.
-TikTok’s audio-visual primacy advantages languages with existing popular media.
-LinkedIn’s professional register expectations advantage languages with established business communication norms.
-Yet these platform-specific language ideologies remain understudied – we lack systematic analysis of how affordances advantage or disadvantage different languages and practices.
+Language ideologies – belief systems concerning language that serve to legitimize and naturalize existing social structures – are necessarily instantiated within technological infrastructures (woolard2020language).
+Contemporary conversational AI systems operationalize Western linguistic ideologies that privilege specific modes of language use, including standardized grammar over vernacular varieties, written over oral forms, explicitness over contextual inference, and monolingualism over multilingual repertoires.
+These ideological commitments inform core design decisions regarding what is classified as high-quality language data, which linguistic features are prioritized for optimization, and which criteria are employed to assess system performance.
+The ideology of a ‘standard language’ constructs socially prestigious varieties as normatively correct while marking other varieties as deficient or erroneous (lippi2012english).
+Large language models (LLMs) are trained predominantly on formal written registers, thereby internalizing and reproducing prescriptive linguistic norms.
+When such models ‘correct’ African American Vernacular English into Standard American English, they operationalize and reinforce raciolinguistic ideologies that frame the former as improper or broken (blodgett2020language).
+Analogous processes occur in multilingual contexts: what is labeled as ‘correct’ Hindi is typically formal written Hindi, which marginalizes and effectively erases Hinglish and diverse regional varieties that constitute the everyday linguistic practices of millions of speakers.
+Platform architectures embed and reproduce particular linguistic assumptions.
+For example, Twitter’s character constraints structurally privilege logographic writing systems relative to alphabetic ones.
+TikTok’s emphasis on audio‑visual content tends to favor languages that already possess extensive, globally circulating popular media repertoires.
+LinkedIn’s normative expectations of a professional register advantage languages with long-established conventions of formal business communication.
+Despite their widespread influence, these platform-specific language ideologies have received limited scholarly attention; systematic analyses are still lacking regarding how platform affordances differentially advantage or disadvantage specific languages and communicative practices.
 3 Methods
 3.1 Data Sources and Coverage
-We analyze comprehensive data on all 7,613 languages documented in the 25th edition of Ethnologue (Lewis et al.,, 2023). Our analysis integrates two primary dimensions:
-Vitality measurement combines:
-  • First-language speaker counts from Ethnologue, ranging from single-digit (nearly extinct languages) to hundreds of millions (Hindi, Arabic, Spanish)
-  • EGIDS (Expanded Graded Intergenerational Disruption Scale) status ratings from 0 (International) to 10 (Extinct), assessing intergenerational transmission and institutional development
-Digitality measurement aggregates presence across:
+We analyze a comprehensive dataset encompassing all 7,613 languages documented in the 25th edition of Ethnologue (eberhard2015ethnologue). The analysis integrates two primary dimensions.
+Vitality is operationalized by combining:
+  • First-language speaker counts from Ethnologue, which range from single-digit values (for languages on the verge of extinction) to several hundred million speakers (e.g., Hindi, Arabic, Spanish)
+  • EGIDS (Expanded Graded Intergenerational Disruption Scale) status ratings, spanning from 0 (International) to 10 (Extinct), which evaluate the degree of intergenerational transmission and the extent of institutional support and development
+Digitality is quantified by aggregating a language’s presence across multiple large-scale digital resources:
   • Common Crawl: 159 billion web pages (November 2023 release)
-  • Wikipedia: 64 million articles across 300+ language editions
-  • Hugging Face: 114,000 datasets and 447,000 models tagged by language
-  • Open Language Archives: 474,000 entries from ELAR, AILLA, DoBeS, and other repositories
-Following Simons et al., (2022), we employ automated language identification using fastText and CLD3 classifiers with validation sampling.
-For each language, we compute normalized scores on each dimension, then apply unsupervised clustering (Gaussian Mixture Models with 3 components) to extract composite vitality and digitality dimensions.
-The representation score is computed as:
-$\text{Representation}=\text{Digitality}_{\text{normalized}}-\text{Vitality}_{\text{normalized}}$ (1)
-Negative scores indicate languages whose vitality exceeds digitality (underrepresented online), positive scores indicate the opposite (overrepresented relative to speaker base).
-3.2 Categorical Classification
+  • Wikipedia: 64 million articles distributed across more than 300 language editions
+  • Hugging Face: 114,000 datasets and 447,000 models annotated with language tags
+  • Open Language Archives Community (OLAC): 474,000 catalogued language resource entries
+Following simons2022assessing, we implement automated language identification using the fastText and CLD3 classifiers, complemented by validation sampling.
+For each language, we calculate normalized prevalence scores across each of the aforementioned corpora and subsequently perform principal component analysis–based dimensionality reduction to derive composite indices of linguistic vitality and digital presence.
+The degree of visibility of a language on digital platforms is quantified by its representation score, which is defined as:
+$\text{Representation}=\text{Digitality}_{\text{normalized}}-\text{Vitality}_{\text{normalized}}\,.$ (1)
+Negative values of this score identify languages whose sociolinguistic vitality exceeds their level of digital presence, indicating underrepresentation in online environments.
+Conversely, positive values correspond to languages that are overrepresented in the digital sphere relative to their speaker base.
 We classify languages into four categories based on their position in vitality-digitality space:
-  • Strongholds (+ Vitality / + Digitality): Both scores above median
-  • Digital Echoes (– Vitality / + Digitality): Below-median vitality, above-median digitality
-  • Fading Voices (– Vitality / – Digitality): Both scores below median
-  • Invisible Giants (+ Vitality / – Digitality): Above-median vitality, below-median digitality
-3.3 Geographic and Colonial Analysis
-To test whether digitality gaps reflect colonial histories, we coded languages by:
-  • Geographic region (Africa, Asia, Europe, Americas, Pacific)
-  • Colonial history (colonizer nation and duration, coded from historical databases)
-  • Official language status (national, regional, none)
-  • Writing system availability and Unicode support
-We employ logistic regression predicting Invisible Giant status from colonial history variables, controlling for speaker population and geographic region.
-3.4 LLM Training Data Analysis
-We analyze language distribution in major LLM training datasets:
-  • The Pile (Gao et al.,, 2021): 800GB English-focused corpus
-  • mC4 (Xue et al.,, 2021): Multilingual C4 covering 101 languages
-  • ROOTS (Laurencçon and the BigScience Workshop,, 2022): BLOOM training data, 46 languages
-  • OSCAR (Suárez et al.,, 2019): Common Crawl extractions, 166 languages
-For each dataset, we quantify token counts by language and correlate with vitality scores to assess whether training data allocation reflects demographic reality or digital bias.
+  • Strongholds (+ Vitality / + Digitality): Entities characterized by both vitality and digitality scores exceeding the sample median.
+  • Digital Echoes (– Vitality / + Digitality): Entities exhibiting below-median vitality combined with above-median digitality.
+  • Fading Voices (– Vitality / – Digitality): Entities for which both vitality and digitality scores fall below the median.
+  • Invisible Giants (+ Vitality / – Digitality): Entities displaying above-median vitality in conjunction with below-median digitality.
 4 Results
+4.1 The Four-Category Framework
+Application of the information visibility framework to all 7,613 documented human languages reveals a pronounced systemic disparity: patterns of linguistic representation within digital ecosystems diverge substantially from patterns of demographic vitality.
+Figure 1 depicts the resulting vitality–digitality matrix, in which each language is positioned according to its real-world demographic strength (x-axis) and its aggregated digital presence across web pages, Wikipedia, datasets, computational models, and archival resources (y-axis).
+The representation score – defined as the difference between digitality and vitality – quantifies this discrepancy, with blue values indicating digital over-representation and red values indicating systematic under-representation.
+Four distinct clusters emerge, each revealing different dynamics of visibility and erasure.
+The geographic spread of these languages can be seen 2, highlighting the stark colonial legacies of invisibility in parts of Africa, Asia, and Oceania.
 Figure 1: Mapping 7613 human languages on vitality (ground presence) and digitality (web presence).
 Figure 2: Geolocating invisible languages
-4.1 The Four-Category Framework
-Applying the invisible information framework to all 7,613 documented human languages reveals a stark reality: linguistic representation in digital ecosystems follows systematic patterns that diverge dramatically from demographic vitality.
-Figure 1 visualizes the vitality-digitality matrix, with each language plotted according to its real-world demographic strength (x-axis) and digital presence across web pages, Wikipedia, datasets, models, and archives (y-axis).
-The representation score – computed as digitality minus vitality – encodes the gap, with blue indicating digital overrepresentation and red indicating systematic underrepresentation.
-Four distinct clusters emerge, each revealing different dynamics of visibility and erasure:
 Figure 3: Strongholds
-Strongholds (+ Vitality / + Digitality): Roughly one-third of languages occupy the upper-right quadrant, exhibiting both demographic robustness and rich digital representation.
-These languages form a plume rising above both zero reference lines, their speaker bases, institutional support, and digital footprints mutually reinforcing.
-This category includes not only global lingua francas but also regionally dominant languages that have successfully translated offline vitality into digital presence.
-From an LLM perspective, these languages already power most training corpora and anchor multilingual benchmarks.
-The policy implication is maintenance rather than intervention – keeping data current while redirecting capacity-building resources toward less fortunate quadrants.
+Strongholds (+ Vitality / + Digitality): Approximately one-third of the languages in the sample fall within the upper-right quadrant, characterized by both high demographic vitality and substantial digital representation (see Figure 3).
+These languages appear as a plume extending above both zero-reference axes, with speaker populations, institutional backing, and online presence functioning as mutually reinforcing factors.
+This group comprises not only global lingua francas but also regionally dominant languages that have effectively converted offline sociolinguistic strength into robust digital visibility.
+From the standpoint of large language models (LLMs), these languages already constitute the bulk of training corpora and serve as primary reference points for multilingual evaluation benchmarks.
+The principal policy implication is thus one of sustained maintenance rather than active remediation, ensuring that available data remain current while reallocating capacity-building efforts and resources to languages situated in more vulnerable quadrants.
 Figure 4: Digital Echoes
-Digital Echoes (– Vitality / + Digitality): Only 6% of languages appear in the upper-left quadrant – languages whose online presence exceeds their dwindling speaker communities.
-Often characterized by historical prestige, liturgical use, or active diaspora networks, these languages demonstrate that digital corpora can outlive real-world vitality.
-They appear as a thin blue veil above the vitality origin, illustrating how documentation and archiving can preserve linguistic material even as living use declines.
-For researchers, Digital Echoes serve as cautionary tales: high digital metrics do not guarantee linguistic health.
-For policymakers, they represent potential leverage points – using strong digital assets to support revitalization before community fluency erodes further.
+Digital Echoes (– Vitality / + Digitality): Only 6% of languages are located in the upper-left quadrant, representing cases in which digital presence exceeds that of shrinking speaker communities (see Figure 4).
+Typically associated with historical prestige, liturgical or ritual functions, and/or active diaspora networks, these languages illustrate that digitally available corpora can persist beyond the decline of face-to-face communicative practices.
+In the visualization, they appear as a thin blue band above the vitality origin, emphasizing how documentation, standardization, and archiving can preserve linguistic material even as inter-generational transmission and everyday use diminish.
+From a research perspective, Digital Echoes underscore a critical methodological limitation: elevated digital indicators (e.g., corpus size, number of online documents, or web traffic) do not equate to robust linguistic health.
+This disconnect arises because most widely used digital metrics are usage-agnostic with respect to (i) the age structure and size of the speaker population, (ii) the domains of actual use (home, school, government, workplace, etc.), (iii) transmission to younger generations, and (iv) speakers’ language attitudes, identity functions, and community control over language resources.
+A language may thus exhibit substantial digital output generated by a small, aging, or highly specialized group (such as clergy, scholars, or heritage enthusiasts), or by nonnative learners and institutions, while the number of fluent, everyday speakers continues to contract.
+Conversely, automated or large-scale digitization (e.g., of historical texts) can inflate digital volume without indicating current communicative vitality or community-driven use.
+For policymakers, these languages constitute strategic leverage points: existing digital infrastructures – including corpora, standardized orthographies, pedagogical materials, and online communities – can be mobilized in support of revitalization policies, such as the development of curricular resources, digital learning tools, and community media, prior to further erosion of any remaining active fluency and functional domains of use.
 Figure 5: Fading Voices
-Fading Voices (– Vitality / – Digitality): The densest cluster, approximately 2,700 languages (36%), hugs the origin and extends into the lower-left.
-These languages suffer a double deficit: small, vulnerable speech communities and vanishingly small digital footprints.
-They form a pale cloud of nearly colorless points, underscoring that when both dimensions are scarce, even the representation gap provides minimal signal.
-From an LLM perspective, they are effectively invisible.
-From a linguistic diversity standpoint, they constitute the front lines of loss – without urgent fieldwork and community-led documentation, these voices will fade beyond empirical recovery.
+Fading Voices (– Vitality / – Digitality): The densest cluster, comprising approximately 2,700 languages (36%), is concentrated near the origin of the space and extends into the lower-left quadrant (see Figure 5).
+These languages exhibit a dual deficit: they are characterized by small, vulnerable speech communities and extremely limited digital presence.
+In the visualization, they appear as a diffuse cloud of nearly colorless points, indicating that when both dimensions are severely constrained, even the representation-gap metric yields minimal informative signal.
+From the perspective of large language models (LLMs), these languages are effectively absent from the training landscape.
+From the perspective of linguistic diversity and language endangerment, however, they represent the bleeding edge of potential loss: without rapid, systematic field documentation and community-driven efforts to record and maintain linguistic practices, these languages are at high risk of disappearing beyond the possibility of empirical reconstruction.
 Figure 6: Invisible Giants
-Invisible Giants (+ Vitality / – Digitality): Approximately 2,000 languages (27%) occupy the lower-right quadrant – the critical focus of our analysis.
-These languages manifest as vivid red points: their vitality extends rightward while digitality hovers below the horizontal reference line, producing the largest positive ”need for action” gap on the color scale.
-These are languages with millions of active speakers but scant digital representation.
-The paradox is profound: robust demographic presence coexists with near-total digital absence.
-The Invisible Giants category reveals the core thesis of this paper: digital exclusion is not simply a function of endangerment or small speaker populations.
-Languages can be vibrant, widely spoken, and intergenerationally transmitted while remaining systematically invisible in the digital ecosystems that increasingly define authoritative knowledge and power contemporary AI systems.
+Invisible Giants (+ Vitality / – Digitality): Approximately 2,000 languages (27%) are located in the lower-right quadrant, which constitutes the primary area of analytical concern in this study (see Figure 6).
+These languages appear as intensely red points: their vitality extends toward the positive end of the horizontal axis, while their digitality remains below the horizontal reference line, yielding the largest positive ”need for action” differential on the color scale.
+These are languages with millions of active speakers but minimal or negligible digital representation.
+The paradox is striking: substantial demographic presence coexists with an almost complete absence from digital domains.
+The Invisible Giants category encapsulates the central argument of this work: digital exclusion cannot be reduced to linguistic endangerment or to small speaker populations.
+Languages may be socio-linguistically robust, widely used, demographically significant, and transmitted across generations, while remaining systematically marginalized or absent within the digital ecosystems that increasingly structure authoritative knowledge production and shape the operation of contemporary AI systems.
 4.2 Theoretical Interpretation: Why Invisible Giants Exist
@@
 This pattern aligns precisely with postcolonial theory’s predictions.
-Said, (1978) established how colonial powers constructed hierarchies positioning non-Western languages as inferior, requiring civilizing intervention.
-Phillipson, (1992) documented how British colonial education policies systematically suppressed vernacular languages, privileging English and select indigenous lingua francas for administrative convenience.
+said1977orientalism established how colonial powers constructed hierarchies positioning non-Western languages as inferior, requiring civilizing intervention.
 These hierarchies did not disappear with formal decolonization – they became embedded in digital infrastructure decisions.
 The infrastructure for digitality – Unicode encoding, keyboard layouts, spell-checkers, search algorithms, content moderation systems – was built primarily by and for Western languages.
-When platforms assess ”notability” for Wikipedia articles, ”quality” for search rankings, or ”representativeness” for training data, they apply standards developed in English-speaking contexts.
-As Noble, (2018) demonstrated for search engines, these ostensibly neutral technical systems encode cultural assumptions that systematically disadvantage marginalized communities.
-The concept of epistemic injustice (Fricker,, 2007) illuminates the mechanism.
+When platforms assess ‘notability’ for Wikipedia articles, ‘quality’ for search rankings, or ‘representativeness’ for training data, they apply standards developed in English-speaking contexts.
+For search engines, these ostensibly neutral technical systems encode cultural assumptions that systematically disadvantage marginalized communities.
+The concept of epistemic injustice sheds light on the mechanism.
 Digital systems commit hermeneutical injustice by failing to provide linguistic resources needed for marginalized communities to interpret and share their experiences.
-When an Invisible Giant language lacks digital presence, speakers cannot access AI tools in their language, cannot find online educational resources, cannot participate in digital commerce on equal footing.
+When an Invisible Giant language lacks digital presence, it is hard for the speakers access AI tools in their language, find online educational resources or just participate in digital commerce on equal footing.
 The system renders their knowledge production illegible, positioning them as passive consumers of content produced in dominant languages rather than active knowledge creators.
@@
 These languages are not ”low-resource” due to inherent deficiency – they are systematically under-resourced by institutions that chose to invest in some languages but not others.
-The framing matters: ”low-resource” naturalizes scarcity, while ”systematically under-resourced” centers the political decisions that produced it (Dotan et al.,, 2023).
+The framing matters: ‘low-resource’ naturalizes scarcity, while ‘systematically under-resourced’ centers the political decisions that produced it (adebara2022towards).
 4.3 Implications for Large Language Models
@@
 This exclusion produces three compounding harms:
-Performance inequality: Models trained predominantly on Stronghold languages perform dramatically worse on Invisible Giants (Joshi et al.,, 2020).
-This performance gap is not merely inconvenient – it determines who can access AI-powered education, commerce, healthcare, and governance.
+  • Performance inequality: Models trained predominantly on Stronghold languages perform dramatically worse on Invisible Giants.
+This performance gap is not merely inconvenient, but also determines who can access AI-powered education, commerce, healthcare, and governance.
 As LLMs become infrastructure, linguistic exclusion becomes infrastructural inequality.
@@
 When training data excludes Invisible Giants, models cannot represent the knowledge systems, historical narratives, and lived experiences of communities speaking those languages.
-As Bender et al., (2021) argue, this constitutes epistemic violence at scale – billions of parameters encoding the message that certain languages and knowledge don’t matter.
+This constitutes epistemic violence at scale – billions of parameters encoding the message that certain languages and knowledge don’t matter.
   • Feedback loops: Poor LLM performance discourages speakers from using their languages digitally, reducing digital content generation, which further reduces training data, which worsens performance.
 This creates self-reinforcing cycles where invisibility begets further invisibility – a digitally mediated language shift mechanism operating at unprecedented scale and speed.
-The geographic map in Figure 2 visualizes these patterns spatially.
-Africa appears predominantly red, indicating languages with higher vitality than digitality.
-Sub-Saharan African languages, despite centuries of oral tradition maintaining rich knowledge systems, remain systematically absent from digital archives (Nekoto et al.,, 2020).
-Meanwhile, Europe appears predominantly blue – even small European languages maintain digital presence exceeding their speaker populations, a legacy of institutional investment in digital infrastructure.
-This is not coincidence but the continuation of colonial-era hierarchies through algorithmic means – what Mohamed et al., (2020) term ”AI colonialism.” Just as colonial powers extracted material resources while positioning colonized peoples as lacking civilization, contemporary AI development extracts training data (often from Global South digital labor) while positioning non-Western languages as ”low-resource” requiring benevolent intervention.
+The geographic map in Figure 2 provides a spatial representation of these patterns.
+Much of Africa appears in red, indicating languages whose sociocultural vitality exceeds their degree of digital representation.
+Many Sub-Saharan African languages, despite being sustained for centuries through robust oral traditions and complex knowledge systems, remain structurally underrepresented or entirely absent in digital repositories.
+By contrast, much of Europe appears in blue, reflecting that even relatively small European languages exhibit a level of digital presence that surpasses what would be expected from their speaker populations alone, largely due to long-term institutional investment in digital infrastructures.
+This distribution does not constitute a random pattern but rather reflects the re-articulation of colonial-era hierarchies through contemporary algorithmic systems.
+Analogous to how colonial regimes extracted material resources while representing colonized populations as lacking civilization, current AI development practices frequently rely on training data derived from digital labor concentrated in the Global South, while simultaneously classifying non-Western languages as ‘low-resource’ and framing them as objects of technocratic or philanthropic intervention.
 4.4 Validation Through Existing Research
-Our empirical findings align with and extend prior fragmented observations.
-Kornai, (2013) established that only 5% of languages will achieve meaningful digital vitality, identifying digital extinction as independent from offline endangerment.
-Our four-category framework provides the theoretical structure Kornai’s work anticipated but did not fully develop: Digital Echoes exemplify his observation that archived languages can be digitally present without living use; Invisible Giants demonstrate his insight that lack of Wikipedia presence predicts inability to digitally ascend.
-Simons et al., (2022) measured digital language support across 143 platforms, finding only 33 languages (0.4%) achieve full capabilities.
-Their methodology focused on technical infrastructure (encoding, fonts, keyboards); our framework adds the vitality dimension, revealing that infrastructure availability does not guarantee usage.
-Some Stronghold languages have infrastructure and active communities; Invisible Giants may have infrastructure but minimal content; Fading Voices lack both.
-The pattern of 27% Invisible Giants validates warnings from Bender et al., (2021) that LLM training data scarcity for most languages is not incidental but structural.
-Their ”stochastic parrots” critique – that LLMs encode biases at scale without understanding – becomes more urgent when we quantify precisely which 27% of demographically robust languages are systematically excluded from training.
-Crucially, our framework reveals what prior work could not: the magnitude of the representation gap.
-It is one thing to observe that some languages lack digital resources; it is another to demonstrate that approximately 2,000 languages with millions of speakers score in the highest ”need for action” category.
-This quantification transforms the problem from anecdotal concern to measurable crisis requiring systematic intervention.
+Our empirical results both corroborate and refine prior, largely fragmentary, observations.
+kornai2013digital estimated that only 5% of the world’s languages are likely to attain substantive digital vitality and argued that digital extinction proceeds largely independently of offline endangerment trajectories.
+The four-category framework we propose supplies the theoretical architecture that Kornai’s analysis implicitly called for.
+In particular, Digital Echoes operationalize his claim that languages preserved in archival resources may exhibit a form of digital presence without corresponding contemporary communicative use, while Invisible Giants instantiate his finding that absence from Wikipedia is a strong predictor of a language’s inability to attain higher levels of digital presence.
+simons2022assessing conducted a large-scale assessment of digital language support across 143 platforms and found that only 33 languages (0.4%) attained full functional capabilities.
+While their methodology primarily emphasizes the dimension of technical infrastructure support, our framework introduces an additional vitality perspective, demonstrating that the mere availability of infrastructure does not ensure actual language use.
+Within our typology, some Stronghold languages exhibit both robust infrastructure and active user communities; Invisible Giants may possess comparable infrastructural provision yet display minimal content production; and Fading Voices are characterized by the absence of both adequate infrastructure and sustained usage.
+The distributional pattern characterized as the 27% Invisible Giants empirically substantiates the concerns articulated by bender2021dangers, namely that the paucity of LLM training data for the majority of the world’s languages is not accidental but structurally embedded in current data collection and curation pipelines.
+The ‘stochastic parrots’ critique – according to which LLMs reproduce and amplify biases at scale without genuine understanding – acquires heightened significance when we quantify which specific 27% of demographically robust languages are systematically marginalized or omitted from training corpora.
+Crucially, our framework sheds light on an aspect that prior work has not quantified: the magnitude of the representation gap.
+It is not only important to document that certain languages lack adequate digital resources, but also to demonstrate that approximately 2000 languages, collectively spoken by millions of individuals, fall into the highest ‘need for action’ category.
+This quantitative characterization reframes the issue from an anecdotal concern into a measurable crisis that necessitates systematic and large-scale intervention.
 5 Discussion
@@
 The systematic concentration of Invisible Giants – 27% of all languages, representing roughly 2,000 languages with millions of speakers yet minimal digital presence – reveals that digital exclusion operates independently of demographic strength.
-This constitutes what we term digital-epistemic injustice: the systematic exclusion of marginalized linguistic communities from AI-mediated knowledge production through denial of both digital infrastructure and epistemic authority.
+This constitutes a kind of digital epistemic injustice: the systematic exclusion of marginalized linguistic communities from AI-mediated knowledge production through denial of both digital infrastructure and epistemic authority.
 This dual exclusion operates at unprecedented scale in LLM development.
-When billions of parameters encode patterns from training data that systematically excludes these 2,000 languages, the resulting models embed the message that certain ways of knowing, certain knowledge systems, and certain communities don’t matter.
+When billions of parameters encode patterns from training data that systematically excludes these 2000 languages, the resulting models embed the message that certain ways of knowing, certain knowledge systems, and certain communities don’t matter.
 The mechanism operates through self-reinforcing feedback loops.
-Initial digital inequality leads to training data scarcity, producing poor LLM performance, discouraging digital language use, reinforcing data scarcity.
+Initial digital inequality leads to training data scarcity, producing poor LLM performance, discouraging digital language use and reinforcing data scarcity.
 Meanwhile, Stronghold languages – already enjoying institutional support and digital infrastructure – attract platform investment, NLP research funding, and commercial applications, widening gaps.
@@
 The concentration of Strongholds in Europe (where even minority languages maintain robust digital presence) while Invisible Giants concentrate in formerly colonized regions reveals historical continuities.
-Colonial language policies positioned indigenous languages as inferior, requiring civilizing intervention; contemporary ”low-resource language” framing positions them as deficient, requiring benevolent development assistance.
+Colonial language policies positioned indigenous languages as inferior, requiring civilizing intervention; contemporary ‘low-resource language’ framing positions them as deficient, requiring benevolent development assistance.
 Both framings naturalize inequality while obscuring its political origins.
-5.2 Theoretical Contributions
-First, we advance epistemic injustice theory by connecting it to digital infrastructure.
-Fricker, (2007)’s framework focused on face-to-face testimonial exchanges and localized hermeneutical resources.
-Helm et al., (2024) extended epistemic injustice to language modeling but treated digitality as given.
-Our analysis reveals that who gets digitized is itself an exercise of power that precedes and enables all subsequent algorithmic processing.
-The decision to invest in Unicode encoding, keyboard development, spell-checkers, and training corpora for some languages but not others determines which communities can practice hermeneutical justice – using linguistic resources to interpret experiences in ways that AI systems recognize as valid.
-When Invisible Giant languages lack these resources, speakers cannot access AI-powered education, healthcare information, or economic opportunities in their languages.
-The system denies them epistemic standing, positioning them as consumers of knowledge produced elsewhere rather than legitimate knowledge creators.
-Second, we demonstrate how postcolonial hierarchies persist in algorithmic systems.
-While Said, (1978) analyzed textual representations and Mohamed et al., (2020) documented extractive labor practices, we show continuities in linguistic hierarchies themselves.
-The vitality-digitality gap reveals how colonial language policies – suppressing vernaculars, imposing European languages, privileging select lingua francas – established hierarchies that LLM training data allocation now reproduces computationally.
-Contemporary AI development exhibits what we term algorithmic colonialism: extracting digital labor from Global South communities while systematically excluding their languages from AI benefits, then marketing AI tools back to those communities as development interventions.
-This mirrors classic colonial resource extraction while adding a distinctly epistemic dimension – not merely extracting material resources but positioning entire knowledge systems as lacking, requiring external validation and technological uplift.
-Third, we problematize the ”low-resource language” framing ubiquitous in NLP.
-This terminology positions underrepresented languages as inherently deficient – they lack resources.
-Our framework reveals that resource scarcity is a political product, not natural fact.
-The 2,000 Invisible Giants are not inherently low-resource; they are systematically under-resourced by institutions that chose to invest in some languages while neglecting others.
-Reframing from ”low-resource” to ”systematically under-resourced” or ”digitally marginalized” centers structural forces rather than naturalizing inequality.
-This shift has implications beyond terminology – it redirects attention from technical data collection challenges to political questions about whose languages receive institutional investment, whose knowledge systems AI encodes, and who benefits from technological development.
 5.2 Limitations and Future Directions
-Methodological limitations: First, our digitality measurement relies on publicly accessible web data, potentially missing private communication in messaging applications.
+Methodological limitations
+First, our digitality measurement relies on publicly accessible web data, potentially missing private communication in messaging applications.
 If Invisible Giant speakers use their languages extensively via WhatsApp, SMS, or encrypted platforms, our analysis underestimates actual digital use.
@@
 Longitudinal tracking of language trajectories across the vitality-digitality space would enable causal inference about what drives languages from one quadrant to another.
-Theoretical extensions: Future research should examine intersections of linguistic marginalization with other forms of oppression.
-How does digital language inequality interact with disability (especially for signed languages’ digital representation)?
+Theoretical extensions
+Future research should examine intersections of linguistic marginalization with other forms of oppression.
+How does digital language inequality interact with special needs (especially for signed languages’ digital representation)?
 With gender, particularly for languages where speaker populations are gendered?
@@
 How do platform affordances systematically advantage or disadvantage specific languages?
-We need granular sociotechnical analysis of how architectural choices embed linguistic assumptions.
+We need granular socio-technical analysis of how architectural choices embed linguistic assumptions.
 Community-centered research methodologies remain rare.
@@
 Developing genuine co-production methodologies where speakers shape research questions, interpret findings, and control data governance represents a methodological frontier essential for decolonizing research practice.
-Empirical extensions: The framework’s generalizability across domains requires testing.
-We developed it for languages; Saurabh Khanna’s broader Invisible Information project applies similar logic to armed conflicts and scholarly research.
-Does the vitality-digitality logic extend to other knowledge domains – traditional medicine, indigenous land management practices, oral historical archives?
+Empirical extensions
+The framework’s generalizability across domains requires testing, since we only developed it for languages, while similar frameworks can extend towards assessing low visibility of certain armed conflicts and scholarly research.
+Does the vitality-digitality logic extend to other knowledge domains – traditional medicine, land management practices, oral historical archives?
 Systematic testing across domains would validate or refine theoretical claims about digital filtering mechanisms.
@@
 Detailed case studies examining language-specific barriers – orthographic debates, institutional politics, diaspora dynamics – would illuminate mechanisms obscured in aggregate analysis.
-Combining computational measurement with ethnographic depth would strengthen explanatory power.
+Combining computational measurement with ethnographic depth would strengthen explanatory power here.
 5.3 Practical Implications
-For AI Developers: Current LLM training treats language representation as a data availability problem – scrape more web pages, digitize more books.
+For AI Developers
+Current LLM training treats language representation as a data availability problem – scrape more web pages, digitize more books.
 Our analysis reveals this approach is insufficient.
@@
 These are prerequisites for content creation, not merely nice-to-have additions.
-Community-controlled data trusts enabling speakers to govern how language data is collected, used, and monetized would shift power dynamics from extractive to collaborative.
+Community controlled data trusts enabling speakers to govern how language data is collected, used, and monetized would shift power dynamics from extractive to collaborative.
 Evaluation metrics must center non-English linguistic features – tone marking accuracy, morphological complexity handling, pragmatic particle usage – rather than English-centric benchmarks like BLEU scores that penalize linguistic differences as errors.
 Performance equity should be measured not just by aggregate metrics but by whether models serve speakers of Invisible Giants as effectively as they serve English speakers.
-For Policymakers: Language planning must integrate digital dimensions.
+For Policymakers
+Language planning must integrate digital dimensions.
 Granting a language official status without digital infrastructure investment produces hollow recognition.
@@
 Just as the Universal Declaration of Linguistic Rights (1996) establishes rights to use one’s language in education and public life, contemporary frameworks must recognize rights to digital language resources, AI services in one’s language, and participation in knowledge production digitized by platforms.
-International funding mechanisms should prioritize Invisible Giants over Fading Voices for digital investment.
-While all endangered languages merit documentation, the 2,000 Invisible Giants offer highest social return on investment – millions of speakers exist who could immediately benefit from digital resources.
-This is not abandoning Fading Voices but strategically allocating limited resources where impact potential is greatest.
-For Researchers: Methodological decolonization requires fundamental shifts in how we conduct research.
+International funding mechanisms should prioritize Invisible Giants for digital investment.
+These offer highest social return on investment – millions of speakers flourish who could immediately benefit from digital resources.
+For Researchers
+Methodological decolonization requires fundamental shifts in how we conduct research.
 Co-designing studies with speaker communities rather than extracting data, compensating language consultants at professional rates, sharing model weights and tools with communities providing training data, publishing in open-access venues with translations into studied languages – these practices should become norms, not exceptions.
@@
 Research excellence should incorporate community endorsement as a criterion, recognizing that scholarship about marginalized communities should serve those communities’ self-determined goals.
-Funding agencies could accelerate these shifts by requiring invisibility impact assessments for digitization grants – explicitly addressing how proposed work will or will not narrow representation gaps – and prioritizing proposals co-designed with affected language communities over researcher-initiated projects.
+Funding agencies could accelerate these shifts by requiring visibility impact assessments for digitization grants – explicitly addressing how proposed work will or will not narrow representation gaps – and prioritizing proposals co-designed with affected language communities over top-down ones.
 5.4 Toward Linguistic Justice in AI
-The Invisible Giants paradox – millions speaking, none listening digitally – crystallizes the crisis.
-Solutions require confronting uncomfortable truths: English dominance in AI is not natural or technologically necessary but reflects historical power structures that continue shaping which languages receive investment, recognition, and computational support.
-Optimistically, concentrated investment could rapidly narrow gaps.
-The methodology exists: measure representation gaps, prioritize Invisible Giants, invest in community-controlled infrastructure, develop multilingual models with performance parity guarantees.
-The economics are feasible – comprehensive digital infrastructure for 2,000 Invisible Giants would cost a fraction of single LLM training runs measuring in hundreds of millions of dollars.
-Pessimistically, current incentives point opposite directions.
-Commercial AI development optimizes for wealthy markets speaking Stronghold languages.
-Academic research chases state-of-the-art performance on English benchmarks, with multilingual work often treated as secondary.
-Platform business models reward engagement metrics that favor content in dominant languages.
-Without structural intervention – regulatory requirements, funding conditionalities, community organizing – market forces will accelerate inequality.
-The choice before us is clear: will AI systems democratize access to information, education, and economic opportunity across linguistic boundaries?
-Or will they accelerate language shift, compressing human diversity toward English-centered homogeneity?
-The answer depends on whether we treat the vitality-digitality gap as a technical resource allocation problem or as the digital-epistemic injustice it truly represents – requiring not just more data, but fundamental redistribution of power, resources, and recognition.
+The Invisible Giants paradox, populations numbering in the millions, whose speakers are effectively unheard in digital spaces – encapsulates the core of the current crisis.
+Addressing this problem requires acknowledging a set of uncomfortable empirical and historical facts: the predominance of English in contemporary AI systems is neither an inherent linguistic inevitability nor a technical necessity.
+Rather, it is a manifestation of enduring geopolitical and socioeconomic power asymmetries that continue to determine which languages receive sustained investment, institutional recognition, and robust computational support.
+From an optimistic standpoint, targeted and adequately resourced interventions could substantially reduce these disparities.
+The requisite methodological toolkit largely already exists: systematically quantify linguistic representation gaps; prioritize structurally marginalized, high-vitality languages; invest in community-governed digital and computational infrastructures; and design multilingual models with explicit guarantees of performance parity or fairness across languages.
+From a realistic view though, prevailing incentive structures are misaligned with these goals.
+Commercial AI development is primarily optimized for high-income markets whose users predominantly speak globally dominant stronghold languages.
+Academic research agendas frequently prioritize incremental improvements on English-centric benchmarks, while multilingual and low-resource language work is often formulated as secondary or peripheral.
+Platform-based business models privilege engagement metrics that systematically favor content in already dominant languages.
+In the absence of deliberate structural interventions – such as regulatory mandates, conditional public funding mechanisms, and sustained community-based organizing – market dynamics are likely to further intensify linguistic and digital inequalities.
+The resulting policy and design choice is stark: Will emerging AI technologies be deployed to democratize access to information, education, and economic opportunity across linguistic boundaries, or will they function as accelerants of language shift, compressing linguistic diversity into an increasingly English-centered monoculture?
+The outcome hinges on whether the gap between linguistic vitality and digital representation is conceptualized merely as a technical resource allocation problem, or as a form of digital and epistemic injustice.
+The latter framing underscores that closing this gap requires not only additional data and infrastructure, but also a more profound redistribution of power, resources, and epistemic recognition across linguistic communities.
 6 Conclusion
-We have demonstrated that linguistic inequality in Large Language Models is not incidental but structural, reflecting continuities from colonial-era hierarchies to contemporary AI development.
-Of the world’s 7,613 languages, roughly 2,000 Invisible Giants with millions of speakers remain digitally underrepresented, concentrating in formerly colonized regions at rates nine times higher than Europe.
-This pattern constitutes digital-epistemic injustice: the systematic exclusion of marginalized linguistic communities from AI-mediated knowledge production through denial of digital infrastructure and epistemic authority.
-As LLMs become infrastructure for education, commerce, and governance, this exclusion threatens to accelerate language shift, compressing human linguistic diversity toward homogenous English-centered multilingualism.
-Yet this outcome is not inevitable.
-Concentrated investment in digital infrastructure for systematically under-resourced languages, community-controlled data governance, and decolonial AI development practices could narrow gaps within a generation.
-The question is one of political will: do we accept AI systems that reproduce colonial hierarchies, or do we demand technology that honors and amplifies the full richness of human linguistic diversity?
-The invisible languages of the LLM universe need not remain invisible.
-Making them visible requires recognizing that their exclusion was never accidental – and that inclusion requires not just more data, but fundamentally reimagining who AI serves and whose knowledge counts.
+We demonstrate that linguistic inequality in Large Language Models (LLMs) is not incidental but structurally embedded, reproducing historical continuities from colonial language hierarchies through to contemporary AI development.
+Of the world’s 7,613 attested languages, approximately 2,000 can be classified as Invisible Giants – languages with millions of speakers that remain substantially underrepresented in digital ecosystems.
+This underrepresentation is disproportionately concentrated in formerly colonized regions, where rates of linguistic exclusion are estimated to be orders of magnitude higher than in Europe.
+We conceptualize this configuration as a form of digital-epistemic injustice: the systematic marginalization of specific linguistic communities within AI-mediated regimes of knowledge production, enacted through constrained access to digital infrastructure and the systematic devaluation of their epistemic contributions.
+As LLMs increasingly function as core socio-technical infrastructure for education, commerce, governance, and cultural production, such exclusion is likely to intensify ongoing processes of language shift.
+This dynamic promotes convergence toward a homogenized, English-dominant model of multilingualism and, in turn, accelerates the erosion of global linguistic diversity.
+This trajectory, however, is not immutable.
+Targeted and sustained investment in digital infrastructure for structurally under-resourced languages, the institutionalization of community-governed data stewardship, and the adoption of explicitly decolonial principles in AI design and governance have the potential to significantly mitigate these disparities within a generational timescale.
+The central question is political rather than purely technical: whether societies will accept AI systems that reproduce and normalize colonial linguistic hierarchies, or instead demand technological infrastructures that recognize, sustain, and amplify the full spectrum of human linguistic diversity.
+The invisible languages within the current LLM ecosystem need not remain invisible.
+Rendering them visible requires acknowledging that their exclusion has been historically and structurally produced rather than accidental, and recognizing that meaningful inclusion entails more than the quantitative expansion of training datasets.
+It necessitates a fundamental reconfiguration of whose interests AI systems are designed to serve and whose knowledge systems are accorded legitimacy and authority.
 Acknowledgments
-We acknowledge research funding from the Fair and Resilient Societies Working Group and the Amsterdam School of Communication Research at the University of Amsterdam.
-We also acknowledge the countless language communities whose knowledge and labor make AI systems possible, yet who rarely benefit from the technologies built on their linguistic resources.
+We acknowledge research funding for this project from the Amsterdam School of Communication Research at the University of Amsterdam.